Farmer has 800 yards of fencing to enclose rectangular garden. Express the area A of the rectangle as a function of the width X of the rectangle. What is the domain of A?
Answer #1
Perimeter = 2L +2w = 800 yds . L=(800-2w) /2 = 400-w . Area = L*W = (400-w) * w= 400w-w^2 . The graph of width vs Area starts at 0, rises to a max, and then drops again to 0 ( this happens when x=0, and when L=0) ( L=0=400-w ,,,,,,,,,w=400) . 0 less than x less than 400,,,,,,,,,,,,, is DOMAIN of Area
answered by: Gianna